# Cross-Cultural Differences in Mental Health Expressions on Social Media

### v3

@@ 6 Background
Depression and anxiety disorders are the most imminent mental health challenges, with the highest contribution to Indian Disability Adjusted Life Years  Sagar et al. (2020). Fear of shame is a primary barrier to mental health recovery in India whereas, for example, substance abuse is the major hurdle in America  Biswas et al. (2016). $71\%$ of Indians exhibited stigma when answering questions about mental health Foundation (2018). Relatedly, somatic symptoms, hypochondriasis, anxiety, and agitation are more commonly seen in Indian patients compared to psychological symptoms than British patients  Gada (1982). While the extent of mental health stigma and treatment (un)availability is often studied, it remains unknown how individuals suffering from mental disorders in India express and seek support on social media.
There is accumulating evidence that suggests language markers of depression vary with demographics such as race Rai et al. (2024); Aguirre and Dredze (2021), immigrant status Mittal et al. (2023a), and geographic location De Choudhury et al. (2017). For instance, Indian and South Africa-based users are less candid in their posts and less likely to exhibit negative emotions in comparison to their Western counterparts De Choudhury et al. (2017). Another study studied the language of Asian users including Indian, Malaysian, and Filipino on Mental Health Support Forum, Talklife Pendse et al. (2019) and found that Indians discuss “wanting or needing friends” more than other countries.

### v4

@@ 6 Background
Depression and anxiety disorders are the most imminent mental health challenges, with the highest contribution to Indian Disability Adjusted Life Years  Sagar et al. (2020). Fear of shame is a primary barrier to mental health recovery in India whereas, for example, substance abuse is the major hurdle in America  Biswas et al. (2016). $71\%$ of Indians exhibited stigma when answering questions about mental health Foundation (2018). Relatedly, somatic symptoms, hypochondriasis, anxiety, and agitation are more commonly seen in Indian patients compared to psychological symptoms  Gada (1982). While the extent of mental health stigma and treatment (un)availability is often studied, it remains unknown how individuals suffering from mental disorders in India express and seek support on social media. There is accumulating evidence that suggests language markers of depression vary with demographics such as race Rai et al. (2024); Aguirre and Dredze (2021), immigrant status Mittal et al. (2023a), and geographic location De Choudhury et al. (2017).